OpenStudy (anonymous):

integral from 0 to 2pi: dx/2+cosx

OpenStudy (anonymous):

\[\int\limits_{0}^{2\Pi} dx/2+cosx\]

OpenStudy (anonymous):

That's....a very long and tedious antiderivative.....If you are allowed to use a calculator, I'd advise just plugging it into the fnint( function on your TI If not, it is a very complicated u substitution.

OpenStudy (anonymous):

isnt dx=2/1+u^2 and cosx=(1-u^2)/(1+u^2) , can we plug it in in terms of u and than solve it

OpenStudy (anonymous):

It looks like the integral will be: = 2 Int ( dx/ ( 2^2 - cos^2x) ) - Int ( cosx dx / ( 2^2 - cos^2x) )